### v2

# Restricted Path Characteristic Function Determines the Law of Stochastic Processes

Abstract.
A central question in rough path theory is characterising the law of stochastic processes. It is established in [I. Chevyrev &\&& T. Lyons, Characteristic functions of measures on geometric rough paths, Ann. Probab. 44 (2016), 4049–4082] that the path characteristic function (PCF), i.e., the expectation of the unitary development of the path, uniquely determines the law of the unparametrised path. We show that PCF restricted to certain subspaces of sparse matrices is sufficient to achieve this goal. The key to our arguments is an explicit algorithm — as opposed to the nonconstructive approach in [I. Chevyrev &\&& T. Lyons, op. cit.] — for determining a generic element $X$ of the tensor algebra ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}=\bigoplus_{n=0}^{\infty}\left(\mathbb{R}^{d}\right)^{\otimes n}$ from its moment generating function. Our only assumption is that $X$ has a nonzero radius of convergence, which relaxes the condition of having an infinite radius of convergence in the literature.
As applications of the above theoretical findings, we propose the restricted path characteristic function distance (RPCFD), a novel distance function for probability measures on the path space that offers enormous advantages for dimension reduction. Its effectiveness is validated via hypothesis testing on fractional Brownian motions, thus demonstrating the potential of RPCFD in generative modeling for synthetic time series generation.
Key words. Path development, path signature, rough path theory, hypothesis testing.
2020 Mathematics Subject Classification:
60L20 (Primary), 62M99, 60G35, 62M07 (Secondary)
$\ast$ Corresponding author.
1. Introduction
Charaterisation of the law of stochastic processes is of enormous significance in stochastic analysis, statistics, and machine learning [16]. The study of faithful and parsimonious representations of laws on the path space, in particular, has a direct impact on algorithm design and efficiency. It has found successful applications in numerical methods of SDE simulations, hypothesis testing on stochastic processes, and generative models for synthetic time series, etc. [25, 24, 23, 26].
The signature of paths, a central concept in rough path theory, can be regarded as non-commutative monomials on the path space. It uniquely determines the underlying path of finite $p$-variations up to a negligible equivalent class; $p\geq 1$. Therefore, characterising the law of stochastic processes is equivalent to characterising the measures on the space of path signatures. It is shown in [7] that the path characteristic function (PCF), i.e., the expectation of the unitary development of the path, determines the law of the signature.
In this paper, we prove that a sparse version of PCF determines the law of the signature under very mild regularity conditions, thus significantly enhancing the results in [7]. Our approach is the natural analogue in the (infinite-dimensional) tensor algebra setting of the classical moment problem, i.e., determination of the law of an $\mathbb{R}$-valued random variable from its moment generating function.
To be more precise, consider
$X=S\big{(}\gamma_{[0,T]}\big{)},$
where $\gamma$ is a stochastic process on $\mathbb{R}^{d}$ and $S$ denotes the signature map. Assume that $\gamma$ is of bounded $p$-variations with $21\leq p<$ 2, so that $X$ takes values of group-like elements in $\bigoplus_{n=0}^{\infty}\left(\mathbb{R}^{d}\right)^{\otimes n}$. The key results in Chevyrev–Lyons [7, Theorem 4.8; Corollary 4.12] ascertain that the characteristic function (see (12) for definition) over $\bigoplus_{k=0}^{\infty}\mathfrak{su}(k)$-valued linear mappings determines the law of $X$, and hence essentially [^1] That is, modulo tree-like equivalence; cf. [10] that of $\gamma_{[0,T]}$. Our paper improves [7] by showing that the same conclusion remains valid when $\bigoplus_{k=0}^{\infty}\mathfrak{su}(k)$ above is replaced by the subalgebra $\bigoplus_{k=0}^{\infty}\mathfrak{so}(k;\mathbb{R})$, or even by the subalgebra of very sparse tridiagonal matrices $\bigoplus_{k=0}^{\infty}\mathfrak{t}(k,\mathbb{R})$ in (1) below. In machine learning applications, this provides enormous benefits to dimension reduction in Cartan development-based distance functions for measures on the path space. See §§5 and 6.
In a much broader context, we study the “generalised moment problem”:
Determine a tensor $X\in\bigoplus_{n=0}^{\infty}\left(\mathbb{R}^{d}\right)^{\otimes n}$ with a nonzero radius of convergence (see §1.4) from its moment generating function.
The only requirement is that the radius of convergence is nonzero, which relaxes the condition of infinite radius of convergence in [7]. We establish an explicit, constructive algorithm for the generalised moment problem in the tensor algebra. The Main Theorem 1.8 states that given the unitary development of an unknown tensor with nonzero radius of convergence, by selecting a few sparse matrices according to certain combinatorial rules, taking the Cartan development of their linear combination, differentiating with respect to the coefficients (“parameters”) $\theta$ in the linear combination, and finally evaluating at $\theta=0$, we are able to recover the tensor.
Our algorithm justifies that the unitary development constitutes a good feature representation of paths in machine learning applications, which will be substantiated in §6. Most notably, the matrices involved in the algorithm are rather sparse: they take values in
$\mathfrak{t}(k,\mathbb{R}):=\bigg{\{}M\in\mathbb{R}^{k\times k}:\,\text{$M+M^{\top}=0$, and the $(i,j)$-entry of $M$}\\ \text{is possibly non-zero only when $|i-j|=1$}\bigg{\}},$ (1)
hence rendering considerable advantages in dimension reduction.
As a corollary, if a tensor $X$ with positive or infinite radius of convergence is nonzero at the $n^{\text{th}}$ level, then for each $k\geq n+1$ one can find a linear map $M:\mathbb{R}^{d}\to\mathfrak{t}(k,\mathbb{R})$ such that the unitary development of $X$ evaluated at $M$ is nonzero. In contrast, [7, Theorem 4.8] establishes that if a tensor $X$ with infinite radius of convergence is nonzero at the $n^{\text{th}}$ level, one can find $M:\mathbb{R}^{d}\to\mathfrak{sp}(m)$ for some $m\geq\max\left\{2,\frac{n}{3}\right\}$ such that the unitary development of $X$ evaluated at $M$ is nonzero. Here $\mathfrak{sp}(m)$, the Lie algebra whose complexification is the symplectic matrix algebra of size $2m\times 2m$, contains $\mathfrak{t}(2m,\mathbb{R})$ as a subspace. The latter result is proved in [7] via polynomial identities in numerical linear algebra works [8, 9] in a non-constructive manner. It only ascertains the existence of such $M$, but gives no clue at all on explicit constructions of $M$.
To describe our results in full rigour, let us first introduce the notions of signature, development, generating function, and radius of convergence in the subsections below.
1.1. Path signature
Consider a path $\gamma:[0,T]\to\mathbb{R}^{d}$ of bounded $p$-variation; $p\in[1,2)$. The signature of $\gamma$, first introduced by K.-T. Chen in 1950’s ([3, 4]), is defined as a sequence of iterated integrals:
$S(\gamma)_{[0,T]}:=\left(1,\gamma^{(1)}_{[0,T]},\ldots,\gamma^{(n)}_{[0,T]},\ldots\right),$
$\gamma^{(n)}_{[0,T]}:=\int_{0<t_{1}<t_{2}<\ldots<t_{n}<T}{\rm d}\gamma_{t_{1}}\otimes\cdots\otimes{\rm d}\gamma_{t_{n}}\,\in\left(\mathbb{R}^{d}\right)^{\otimes n}$ (2)
that takes values in the tensor algebra:
${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}:=\bigoplus_{n=0}^{\infty}\left(\mathbb{R}^{d}\right)^{\otimes n},$ (3)
with the convention $\left(\mathbb{R}^{d}\right)^{\otimes 0}\equiv\mathbb{R}$. Here the iterated integral in (2) is understood in Young’s sense [17]. It is easy to see that the signature is invariant under time-reparametrisation; hence, we shall drop the subscript [0,T] from $S(\gamma)$ from now on.
The mapping $\gamma\mapsto S(\gamma)$ defines a group homomorphism from the path group (with concatenation $\ast$ and path-inverse $\overleftarrow{\bullet}$ as group operation and inverse) into ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$, such that
$S(\gamma\ast\eta)=S(\gamma)\otimes S(\eta)\qquad\text{and}\qquad S(\gamma)\otimes S\left(\overleftarrow{\gamma}\right)={\bf 1}:=(1,0,0,\ldots).$ (4)
The first identity in (4), first established in [5], is known as Chen’s identity.
The signature contains all essential information of a path. More precisely, Chen showed in [5] that two irreducible piecewise $C^{1}$-paths cannot have the same signature unless they differ only by a translation or reparametrisation. This result was extended to BV-paths by Hambly–Lyons in [10], proving that the signature uniquely determines BV-paths modulo tree-like equivalence.
Besides its theoretical importance, as a mathematically principled feature that enjoys desirable analytic-geometric properties (e.g., universality and characteristicity), the signature establishes itself as a powerful tool in machine learning applications, especially for time series data modelling. It serves as an efficient feature extractor of time series and streaming data and, when integrated with appropriate machine learning techniques, signature-based approaches have shown success in quantitative finance, early detection of disease, and online character recognition, etc. [25, 24, 23, 26]. See also Lyons’ ICM talk [16] and the recent works [6, 12, 15].
1.2. Path development
An important tool for the study of signature is the (Cartan) development of a path. Consider a Lie group $G$ with Lie algebra ${\mathfrak{g}}$. The development of a path $\gamma:[0,T]\to\mathbb{R}^{d}$ into $G$ is a map
${\mathscr{D}}_{\gamma}:{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)\longrightarrow G$
given (formally) by
${\mathscr{D}}_{\gamma}(M):=I_{k}+\sum_{n=1}^{\infty}\int_{0<t_{1}<t_{2}<\ldots<t_{n}<T}M\left({\rm d}\gamma_{t_{1}}\right)\bullet\cdots\bullet M\left({\rm d}\gamma_{t_{n}}\right),$ (5)
where $I_{k}$ is the identity of $G$. This is well defined when $\gamma$ is a path of bounded $p$-variation; $21\leq p<$ 2. The symbol $\bullet$ denotes matrix multiplication, and ${\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)$ is the space of linear transforms from $\mathbb{R}^{d}$ to ${\mathfrak{g}}$. The formula (5) is closely related to the signature of $\gamma$ defined in (2).
The path development plays a central role in the study of signature inversion [19] and the uniqueness of expected signature; cf. e.g., [7, 10, 2]. Chevyrev–Lyons [7] proved that, for suitably chosen ${\mathfrak{g}}$, the development is a universal, characteristic feature. Moreover, the development takes values in Lie groups of finite dimensions (independent of the path dimension $d$), thus serving as a promising tool in machine learning applications for overcoming the “curse of dimensionality” inherent to the path signature. See [15] and the references cited therein.
For $G=U(k)=$ the unitary group and ${\mathfrak{g}}={\mathfrak{su}}(k)=$ the unitary algebra (i.e., the space of skew-Hermitian matrices $M$ with $M+M^{*}=0$), we call ${\mathscr{D}}_{\gamma}$ the unitary development of $\gamma$. In machine learning papers (e.g., [15, 14]), it is also known as the “unitary feature”.
1.3. Generating function of a tensor
Let $x\equiv\{x_{n}\}_{n=0}^{\infty}$ be a real sequence. Its generating function is given by the (formal) complex series:
$\phi_{x}:\mathbb{C}\longrightarrow\mathbb{C},\qquad\phi_{x}(\lambda)=\sum_{n=0}^{\infty}x_{n}\lambda^{n}.$ (6)
For $\lambda\in\mathbb{C}$ with $|\lambda|<\rho(x):=$ the radius of convergence of the above power series, we may differentiate the series term-by-term. Thus, with $\phi_{x}$ given, we may recover $x$ via
$x_{n}=\frac{1}{n!}\frac{{\rm d}^{n}}{{\rm d}\lambda^{n}}\phi_{x}(\lambda)\bigg{|}_{\lambda=0}.$ (7)
Now let us generalise the above construction to tensors.
Definition 1.1.
For each given $k=1,2,3,\ldots$ and tensor $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$, the generating function of $X$ of order $k$ is the mapping
$\Phi_{X}^{(k)}:{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)\,\longrightarrow\,{\mathfrak{gl}(k,\mathbb{C})},$
$\Phi_{X}^{(k)}(M):={\widetilde{M}}(X).$ (8)
Here ${\mathfrak{gl}(k,\mathbb{C})}\cong\mathbb{C}^{k\times k}$ is the algebra of $k\times k$ complex matrices, and ${\widetilde{M}}:{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}\to{\mathfrak{gl}(k,\mathbb{C})}$ is the canonical extension of $M$ defined as follows:
${\widetilde{M}}$ is the algebra homomorphism determined by ${\widetilde{M}}(1)=I_{k}$ and that
${\widetilde{M}}\left(e_{i_{1}}\otimes\cdots\otimes e_{i_{\ell}}\right):=M\left(e_{i_{1}}\right)\bullet\cdots\bullet M\left(e_{i_{\ell}}\right)\text{ for any }\ell\in\mathbb{N}\text{ and }i_{1},\ldots,i_{\ell}\in\{1,\ldots,d\}.$ (9)
We also denote the generating function of $X$ as
$\Phi_{X}:\bigoplus_{k=0}^{\infty}{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)\,\longrightarrow\,\bigoplus_{k=0}^{\infty}{\mathfrak{gl}(k,\mathbb{C})},$
$\Phi_{X}\Big{|}_{{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)}=\Phi_{X}^{(k)},$
which respects the natural grading in $k\in\mathbb{N}$.
We define ${\widetilde{M}}$ by (1.1) since
$\{1\}\cup\big{\{}e_{i_{1}}\otimes\cdots\otimes e_{i_{\ell}}:\ell\in\mathbb{N}\text{ and }i_{1},\ldots,i_{\ell}\in\{1,\ldots,d\}\big{\}}$
forms a Hamel basis for ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$. Notice too that the defining equation (1.1) for the generating function is formal — consider the projection (a.k.a. truncation) of a tensor onto the $n^{\text{th}}$ level:
$\pi_{n}:{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}\to\left(\mathbb{R}^{d}\right)^{\otimes n}.$ (10)
In view of (1.1), the action of ${\widetilde{M}}$ on $\pi_{n}(X)$ involves a product of $n$ matrices. To make (1.1) rigorous, we require that each $n$-fold product in (1.1) converges in ${\mathfrak{gl}(k,\mathbb{C})}$ as the tensor degree $n\to\infty$. To fix the idea, in this paper we always adopt the Hilbert–Schmidt norm on ${\mathfrak{gl}(k,\mathbb{C})}$.
Definition 1.1 of the generating function can be extended to any Lie algebra $\mathfrak{g}$ and normed unital algebra $\mathfrak{a}$. One defines $\Phi_{X}:\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{g}\right)\to\mathfrak{a}$ again via $\Phi_{X}(M):={\widetilde{M}}(X)$, where ${\widetilde{M}}$ is given by (1.1) with $\bullet$ being the multiplication in $\mathfrak{a}$. If $X\in\mathcal{G}\left(\mathbb{R}^{d}\right)$, then $\Phi_{X}$ takes values in $G$, the Lie group of $\mathfrak{g}$. On the other hand, if $X$ is in ${\rm Lie}\left[\mathcal{G}\left(\mathbb{R}^{d}\right)\right]$, the Lie algebra of $\mathcal{G}\left(\mathbb{R}^{d}\right)$, then $\Phi(X)\in\mathfrak{g}$ [^2] The Lie algebra of $\mathcal{G}\left(\mathbb{R}^{d}\right)$ is denoted as $\mathcal{L}((\mathbb{R}^{d}))$ in [17]. We avoid using this notation and reserve $\mathcal{L}$ for the space of linear maps. Here and hereafter,
$\mathcal{G}\left(\mathbb{R}^{d}\right):=\left\{\text{group-like elements in ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$}\right\},$
whose Lie algebra ${\rm Lie}\left[\mathcal{G}\left(\mathbb{R}^{d}\right)\right]$ is spanned by $\left[e_{i_{1}},\left[e_{i_{2}},\cdots,\left[e_{i_{\ell-1}},e_{i_{\ell}}\right]\right]\right]$ where $\{e_{1},\ldots,e_{d}\}$ is a basis for $\mathbb{R}^{d}$. See [17, §2.2.4].
Example 1.2.
When $d=1$ and $k=1$, the 1-dimensional matrix algebra ${\mathfrak{su}}(1)$ equals ${\sqrt{-1}}\mathbb{R}$. We view $\mathbb{R}^{\otimes n}\cong\mathbb{R}$ and hence the tensor algebra satisfies
${\mathcal{T}}(\mathbb{R})\cong\bigoplus_{n=0}^{\infty}\mathbb{R}=\text{the space of real sequences}.$
Thus, for $x\in{\mathcal{T}}(\mathbb{R})$ one identifies $\pi_{n}(x)=x_{n}1^{\otimes n}\in\mathbb{R}^{\otimes n}$ and
$x=(x_{0},x_{1},x_{2},\ldots)\sim\sum_{n=0}^{\infty}x_{n}1^{\otimes n}.$ (11)
In this way, for each $\lambda\in{\mathcal{L}}(\mathbb{R};{\sqrt{-1}}\mathbb{R})$ we view
$\widetilde{\lambda}\left(x_{n}1^{\otimes n}\right)=x_{n}\overbrace{\lambda(1)\cdots\lambda(1)}^{n\text{ times }}\equiv x_{n}\lambda^{n},$
with the obvious abuse of notation $\lambda(1)\equiv\lambda$.
Therefore, in view of the series notation in (11), we may express $\phi_{x}(\lambda)=\sum_{n=0}^{\infty}x_{n}\lambda^{n}=\Phi_{x}(\lambda)$. That is, the generating function for tensors over $\mathbb{R}^{1}$ into 1-dimensional unitary matrix algebra agrees with the classical notation in (6) for sequences. This also explains why we need to consider complex series for (6).
Observe the structural resemblance of the integrand of the development in (5) with the canonical extension in (1.1) of the generating function. Indeed, we have the following result.
Proposition 1.3.
The development and signature of a BV-path $\gamma:[0,T]\to\mathbb{R}^{d}$ are related by the following: for any $k=1,2,3,\ldots$ and $M\in{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)$, it holds that
${\mathscr{D}}_{\gamma}(M)={\widetilde{M}}\big{(}S(\gamma)\big{)}.$
Proof.
See [14, Appendix A]; ${\mathfrak{su}}(k)$ can be replaced by any finite-dimensional Lie algebra. ∎
In passing, we remark that Definition 1.1 for the generating function is closely related to the notion of the characteristic function $\phi_{X}$ for a $G$-valued random variable $X$, where $G$ is a topological group (see Heyer [11] and Chevyrev–Lyons [7, §1]):
$\phi_{X}(M):=\mathbb{E}\left[{\widetilde{M}}(X)\right]\qquad\text{where $M$ isany unitary representation of $G$.}$ (12)
In light of Proposition 1.3, if $X=S(\gamma)$ for some random BV-path $\gamma$, then $\phi_{X}(M)$ equals the expectation of the development ${\mathscr{D}}_{\gamma}(M)$. Here $G$ can be taken as $\mathcal{G}\left(\mathbb{R}^{d}\right)$, the group-like elements in the tensor algebra ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$; cf. [7, §3].
1.4. The generalised moment problem; radius of convergence
One central question concerning the (unitary) development is whether the analogue of (7) holds in the case of tensors.
Question 1.4.
Given the mapping $\Phi:\bigoplus_{k=0}^{\infty}{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)\to\bigoplus_{k=0}^{\infty}{\mathfrak{gl}(k,\mathbb{C})}$ graded by $k\in\mathbb{N}$, which is known to be the generating function of some $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$, can one solve for $X$?
We shall refer to it as the generalised moment problem [^3] Indeed, in Chevyrev–Lyons [7, §6.1] “the moment problem” refers to our generalised moment problem specialised to the case of expected signatures. for $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$.
Of particular interests are the following choices of $X$:
  • Signature — $X=S(\gamma)$ for some path $\gamma$, e.g., of bounded $p$-variation for $21\leq p<$ 2;
  • Expected signature — $X=\mathbb{E}[S(\gamma)]$ for some random path $\gamma$.
By virtue of Proposition 1.3, in the above cases Question 1.4 specialises to the following
Question 1.5.
Does the (expected, resp.) development uniquely determines the (expected, resp.) signature of a path?
Notice that certain conditions must be imposed on $X$ to get an affirmative answer to Question 1.4. Even in the simplest case of real sequences (Example 1.2), to solve for each $x_{n}$ via (7), the radius of convergence of the power series in (6) must be non-zero.
[…]
Definition 1.6.
The radius of convergence of a tensor $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$, written as $\rho(X)$, is the radius of convergence of the (complex) power series
$\lambda\longmapsto\sum_{n=0}^{\infty}\left\|\pi_{n}(X)\right\|\lambda^{n}.$
Note that $\rho(X)\in[0,\infty]$. For any $R\geq 0$, we further set
${\mathscr{E}}_{R}:=\left\{X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}:\,\rho(X)\geq R\right\},\qquad\widehat{{\mathscr{E}}}:=\bigcup_{R>0}{\mathscr{E}}_{R},\qquad\text{ and }{\mathscr{E}}_{\infty}:=\bigcap_{R>0}{\mathscr{E}}_{R}.$
The radius of convergence $\rho(X)$ of a tensor $X$ is closely related to its generating function $\Phi_{X}$ as in Definition 1.1, as shown by the following
Proposition 1.7.
Given $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$ and $k=1,2,3,\ldots$.
  1. (1)
    If $X\in{\mathscr{E}}_{\infty}$ (i.e., $\rho(X)=\infty$), then the generating function $\Phi_{X}$ in (1.1) is well defined on the domain ${\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)$.
  2. (2)
    For each $R>0$, if $X\in{\mathscr{E}}_{R}$ (i.e., $\rho(X)\geq R$), then $\Phi_{X}$ is well defined on the subdomain $\left\{M\in{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right):\,\text{the operator norm of $M$}<R\right\}$.
  3. (3)
    $X\in{\mathscr{E}}_{\infty}$ if and only if it lies in the closure of ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$ with respect to the coarsest topology for which the following holds: for any normed algebra $\mathcal{A}$ and any algebra homomorphism $M\in{\rm Hom}\left(\mathbb{R}^{d},\mathcal{A}\right)$, the canonical extension ${\widetilde{M}}:{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}\to\mathcal{A}$ is continuous.
Proof.
Clearly $(2)\Rightarrow(1)$. To see (2), we simply note that $\|\widetilde{M}(\pi_{n}(X))\|\leq\|\pi_{n}(X)\|\cdot\|M\|_{\rm op}^{n}$ for each $n\in\mathbb{N}$, where $\|M\|_{\rm op}$ is the operator norm of $M$. Finally, (3) follows from [7, Corollary 2.5], with $\Psi$ chosen to be the singleton of the $\ell_{1}$-norm on $\mathbb{R}^{d}$ therein. ∎
1.5. Main Theorem
Our main result of this paper is to give an affirmative answer to Question 1.4, under the assumption $\rho(X)>0$ — equivalently, $X\in\widehat{{\mathscr{E}}}$. In fact, we shall obtain an explicit algorithm to solve for $X$ from the generating function $\Phi$:
Theorem 1.8.
Given a mapping
$\Phi:\bigoplus_{k=0}^{\infty}{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)\to\bigoplus_{k=0}^{\infty}{\mathfrak{gl}(k,\mathbb{C})}$
graded by $k\in\mathbb{N}$. Suppose that $\Phi$ is the generating function of some $X\in\widehat{{\mathscr{E}}}$, i.e., a tensor with a nonzero radius of convergence. One can determine $X$ explicitly by the following procedure:
(i) Write
$X=\sum X^{W}e_{w_{1}}\otimes\cdots\otimes e_{w_{n}},\qquad W\equiv\left(w_{1},\ldots,w_{n}\right),$
where $\{e_{1},\ldots,e_{d}\}$ is the Cartesian basis for $\mathbb{R}^{d}$ and the summation is taken over all possible $n\in\mathbb{N}$ and ordered $n$-tuples $W\in{\{1,\ldots,d\}}^{n}$. It suffices to determine $X^{W}$ for each $W$.
(ii) To this end, fix any $k\geq n+1$ and take the linear mappings
${{\mathcal{M}}[W]}_{1},\ldots,{{\mathcal{M}}[W]}_{d}\in{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{so}}(k,\mathbb{R})\right)$
defined for each $i,j\in\{1,\ldots,d\}$ as follows:
${{\mathcal{M}}[W]}_{i}(e_{j}):=\begin{cases}\sum^{\prime}_{t}\left(E^{p^{W}(i|t)}_{p^{W}(i|t)+1}-E^{p^{W}(i|t)+1}_{p^{W}(i|t)}\right)\quad\text{ if }i=j\in\{w_{1},\ldots,w_{n}\},\\ 0\qquad\text{ if }i\notin\{w_{1},\ldots,w_{n}\}\text{ or }i\neq j.\end{cases}$
Here, $p^{W}(i|t)$ is the location where $i$ appears for the $t^{\text{th}}$ time in $W=(w_{1},\ldots,w_{n})$, and $\sum^{\prime}_{t}$ means summation over $t$ from $1$ up to the total number of times that $i$ appears in $W$. The symbol $E^{a}_{b}$ designates the $k\times k$ matrix whose $(a,b)$-entry is $1$ and all the other entries are 00.
(iii) Form the linear combination
${{{\mathcal{M}}[W]}_{(\theta)}}=\sum_{i=1}^{d}\theta_{i}{{\mathcal{M}}[W]}_{i},$
where $\theta=(\theta_{1},\ldots,\theta_{d})^{\top}\in\mathbb{R}^{d}$ are “parameters”. Then
$X^{W}=\frac{1}{{\mathcal{C}_{W}}}\times\left\{\text{the $(1,n+1)$-entry of thematrix }\frac{\partial^{n}\Phi\left({{{\mathcal{M}}[W]}_{(\theta)}}\right)}{\partial\theta_{w_{1}}\cdots\partial\theta_{w_{n}}}\Bigg{|}_{\theta=0}\right\},$ (13)
where the combinatorial constant ${\mathcal{C}_{W}}$ equals the product over each $i\in{\{1,\ldots,d\}}$ of the factorial of the number of times that $i$ appears in $W$.
Despite the lengthiness of the statement of Theorem 1.8, we emphasise its explicit and constructive nature. Every input (i.e., constant ${\mathcal{C}_{W}}$ and matrices ${{{\mathcal{M}}[W]}_{(\theta)}}$) in the formula (13) depends only on the word $W$ and is purely combinatorial. Detailed explanations for the notations can be found in §2 below.
An immediate consequence of Theorem 1.8 is the characteristic property of the generating function. That is, generating functions (viewed as the map $X\mapsto\Phi_{X}$) separate points over $\widehat{{\mathscr{E}}}$.
Corollary 1.9.
For any $X\in\widehat{{\mathscr{E}}}$ with $\pi_{n}(X)\neq 0$ and any $k\geq n+1$, we can find $M\in{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$ such that $\Phi_{X}(M)\neq 0$. Here $\Phi_{X}$ is the generating function of $X$ as in Definition 1.1, and $\mathfrak{t}(k,\mathbb{R})$ is the space of tridiagonal antisymmetric $k\times k$ matrices as in Equation 1.
This result should be compared with Chevyrev–Lyons [7, Theorem 4.8], which ascertains that for any $X\in{\mathscr{E}}^{\infty}$ with $\pi_{n}(X)\neq 0$, we can find $M\in{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{sp}(m)\right)$ for some $m\geq\max\left\{2,\frac{n}{3}\right\}$ such that $\Phi_{X}(M)\neq 0$. Here $\mathfrak{sp}(m)$ is the Lie algebra of (the real form of the) symplectic group of $2m\times 2m$ real matrices, which contains $\mathfrak{t}(2m,\mathbb{R})$ as a subset. The proof in [7] is based on the polynomial identities in the numerical linear algebra literature [8, 9].
Given the above discussions on radius of convergence, we can say more about Question 1.5.
  • As is shown in Chevyrev–Lyons [7, Corollary 2.5 and §5], the signature of a $p$-rough path for any $p\geq 1$ has an infinite radius of convergence. Hence, its moment generating function uniquely determines the signature. In particular, for $p\in[1,2)$ the signature can be defined as in (2) in the sense of Young’s integrals, and the signature of any geometric rough path in $\mathbb{R}^{d}$ lies in $\mathcal{G}\left(\mathbb{R}^{d}\right)$, the set of group-like elements in tensor algebra. In this case, the uniqueness of unitary development holds.
  • The case for the expected signature is more complicated. The infinity of radius of convergence for the expected signature has been verified for several popular stochastic processes on a fixed time horizon, e.g., fractional Brownian motions with Hurst parameter in $[\frac{1}{2},1)$ ([21]), as well as certain Gaussian and Markovian rough paths ([7, Examples 6.7 and 6.8]). It is, nonetheless, considerably challenging to study processes up to a random time. The only known results in this regard, to the best of our knowledge, pertain to the Brownian motion up to the first exit time from a bounded domain $\Omega\subset\mathbb{R}^{d}$ (a.k.a. the “stopped Brownian motion”). Using PDE approaches pioneered by Lyons–Ni [18], it has been established in [1, 13] that the expected signature of stopped Brownian motions on $C^{2,\alpha}$-domains in $\mathbb{R}^{d}$ for $2\leq d\leq 8$ has a finite radius of convergence.
1.6. Applications to numerical analysis and machine learning
The use of signature and path development in machine learning is on the rise, due to the characteristic and universal properties. The signature serves as an efficient feature extractor of time series and streaming data. When integrated with appropriate machine learning techniques (e.g., deep learning and gradient boosting trees), signature-based approaches have shown success across various fields — quantitative finance, early detection of disease, and online character recognition, etc. [25, 24, 23, 26].
However, the signature is prone to the curse of dimensionality, especially in tasks involving high-dimensional time series. To this end, [14] proposed the path development layer, which is data-adaptive and significantly reduces dimensions while preserving essential information of the signature feature. Built on top of the expected unitary development, [15] further proposed the so-called path characteristic function (PCF), which proves to be a general metric on the space of distributions on the unparameterised path space. Successful applications of PCF have been found in hypothesis testing on time series distributions and in synthetic time series generation.
The theoretical justification of the PCF distance is essentially based upon the characteristic property of the (expected) unitary development. In Theorem 1.8 and its proof, we specify a reduced set of sparse linear maps in $\mathcal{L}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)$, under which the corresponding unitary development uniquely determines the unparameterised path. It motivates us to consider the “restricted PCF distance”, which characterises the law of time series while reducing the computation complexity compared with the PCF distance in [15]. This will be introduced in §6, with demonstration of its efficacy in hypothesis testing on the fractional Brownian motion.
1.7. Organisation
The remaining parts of the paper are organised as follows.
In §2, we introduce some notations that shall be used throughout this paper. Then, in §3, a combinatorial identity on derivatives of the unitary development of special matrices is presented. Our main Theorem 1.8 is proved in §4. The following section §5 is devoted to the special case of the (expected) signature. Next, in §6, we investigate the application of our main Theorem 1.8 to hypothesis testing, via the concept of restricted path characteristic function distance introduced therein. Some concluding remarks are given in §7. Appendix A presents the details of the numerical experiments on hypothesis testing.
2. Notations and Nomenclature
In this preliminary section, we introduce some notations that will be used throughout. We assume that all paths in consideration take values in $\mathbb{R}^{d}$, with $d$ fixed once and for all.
2.1. Words
We denote
${\bf Word}(n):=\Big{\{}W:\,\text{$W$ is a word of length $n$ fromthe alphabet $\{1,\ldots,d\}$}\Big{\}}.$
That is, $W\in{\bf Word}(n)$ is an ordered $n$-tuple $W=\left(w_{1},\ldots,w_{n}\right)$ where $w_{i}\in{\{1,\ldots,d\}}$ for each $i\in{\{1,\ldots,n\}}$. As a set ${\bf Word}(n)={\{1,\ldots,d\}}^{n}$; each $w_{i}$ appearing in $W$ is called a letter of $W$.
A generic tensor $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$ can be expressed as the (formal) sum:
$X=\sum_{n=0}^{\infty}\,\,\sum_{W\in{\bf Word}(n)}X^{W}e_{W}$ (14)
where, for each $W=\left(w_{1},\ldots,w_{n}\right)$ as above, $X^{W}\in\mathbb{R}$ is a real coefficient and $e_{W}\equiv e_{w_{1}}\otimes\cdots\otimes e_{w_{n}}$.
From an alternative perspective, we can view each $W\in{\bf Word}(n)$ as a multi-index with valency $|W|=n$. One may take partial derivatives via
$\frac{\partial^{n}}{\partial x^{W}}=\frac{\partial^{n}}{\partial x^{w_{1}}\cdots\partial x^{w_{n}}},$
with possible repetitions in $\{w_{1},\ldots,w_{n}\}$.
Denote by $\Sigma_{n}$ the permutation group of $n$ letters, i.e., the set of bijections from ${\{1,\ldots,n\}}$ to itself. One has the natural group action
$\Sigma_{n}\times{\bf Word}(n)\rightarrow{\bf Word}(n)$
given by
$\sigma\cdot\left(w_{1},\ldots,w_{n}\right):=\left(w_{\sigma(1)},\ldots,w_{\sigma(n)}\right).$
One may define an equivalence relation on ${\bf Word}(n)$:
$I\sim J\qquad\text{if and only if}\qquad I=\sigma(J)\text{ for some }\sigma\in\Sigma_{n}.$
That is, $I\sim J$ means that they differ by a permutation, hence have the same length and same (unordered) indices/letters.
2.2. Multiplicity and position functions
For each word $W\in{\bf Word}(n)$ and each letter of it, we define the multiplicity and position functions:
Definition 2.1.
Given $W\in{\bf Word}(n)$ where $n=1,2,3,\ldots$.
  1. (1)
    For $i\in{\{1,\ldots,d\}}$, we set $r^{W}(i):=$ the multiplicity of $i$ in $W$, i.e., the number of times that $i$ appears in $W$.
  2. (2)
    For $i\in{\{1,\ldots,d\}}$ and $t\in\left\{1,\ldots,r^{W}(i)\right\}$, we set $p^{W}(i|t):=$ the position that $i$ appears in $W$ for the $t^{\text{th}}$ time. We put $p^{W}(i|t)=\emptyset$ if $r^{W}(i)=0$.
Example 2.2.
Consider the case $d=4$, $n=5$, and $W=(12221)$. Then
$r^{W}(1)=2,\quad r^{W}(2)=3,\quad r^{W}(3)=r^{W}(4)=0;$
as well as
$p^{W}(1|1)=1,\quad p^{W}(1|2)=5,\quad p^{W}(2|1)=2,\quad p^{W}(2|2)=3,\quad p^{W}(2|3)=4,$
$\qquad\text{and }\quad p^{W}(i|t)=\emptyset\quad\text{ for all the other $i\in\{1,2,3,4\}$ and $t\in\left\{1,\ldots,r^{W}(i)\right\}$}.$
Note that for any $W\in{\bf Word}(n)$, $i\in{\{1,\ldots,d\}}$, and $t\in\left\{1,\ldots,r^{W}(i)\right\}$, we have
$r^{W}(i)\in\{0,1,\ldots,n\}\quad\text{and}\quad p^{W}(i|t)\in{\{1,\ldots,n\}}.$
The function $t\mapsto p^{W}(i|t)$ is strictly increasing for fixed $i$. One also has the simple identity:
$\sum_{i=1}^{n}r^{W}(i)=n\qquad\text{ for any }W\in{\bf Word}(n).$
2.3. Matrices
Fix $k\in\mathbb{N}$ and assume that all the matrices in this paragraph are of size $k\times k$. We use the symbol $E^{a}_{b}$ ($a,b\in{\{1,\ldots,k\}}$) to denote the elementary matrices:
$E^{a}_{b}\in{\mathfrak{gl}(k,\mathbb{C})}\quad\text{such that the $(a^{\prime},b^{\prime})$-entry is $\delta_{a}^{a^{\prime}}\delta^{b}_{b^{\prime}}$ for any $a^{\prime},b^{\prime}\in{\{1,\ldots,k\}}$}.$
That is, $E^{a}_{b}$ is the matrix whose $(a,b)$-entry is $1$ and all the other entries are 00.
The following identity will be used repetitively:
$E^{a}_{b}\bullet E^{p}_{q}=\delta_{b}^{p}E^{a}_{q}\qquad\text{for any }a,b,p,q\in{\{1,\ldots,k\}}.$ (15)
Throughout we use $\bullet$ to denote the matrix multiplication. Notice also that
[…]
Here ${\mathfrak{so}}(k,\mathbb{R})\subset{\mathfrak{su}}(k)$ is the Lie algebra of real skew-symmetric $k\times k$ matrices, which has dimension $\frac{k(k-1)}{2}$. The matrices ${{\mathcal{M}}[W]}_{1},\ldots,{{\mathcal{M}}[W]}_{d}$ in Theorem 1.8 are central to our constructions. Here they are chosen to take values in ${\mathfrak{so}}(k,\mathbb{R})$, but see also Remark 4.1.
3. Derivatives of the development with respect to hyperparameters
Here and hereafter, we take a tensor $X\in\widehat{{\mathscr{E}}}$.
Define for a hyperparameter $\theta=(\theta_{1},\ldots,\theta_{d})\in\mathbb{R}^{d}$ the matrix
${M_{(\theta)}}:=\sum_{i=1}^{d}\theta_{i}M_{i},$ (16)
where $M_{1},\ldots,M_{d}$ are arbitrary linear maps in ${\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)$. The generating function of $X$ acting on ${M_{(\theta)}}$ can be computed as follows:
$\Phi_{X}({M_{(\theta)}})$ $:=\widetilde{\sum_{i=1}^{d}\theta_{i}M_{i}}\,\left(\sum_{n=0}^{\infty}\,\,\sum_{W\in{\bf Word}(n)}X^{W}e_{W}\right)$
$=\sum_{n=0}^{\infty}\,\,\sum_{W=\left(w_{1},\ldots,w_{n}\right)\in{\bf Word}(n)}X^{W}\left\{\sum_{i_{1}=1}^{d}\theta_{i_{1}}M_{i_{1}}\left(e_{w_{1}}\right)\right\}\bullet\cdots\bullet\left\{\sum_{i_{n}=1}^{d}\theta_{i_{n}}M_{i_{n}}\left(e_{w_{n}}\right)\right\}$
$=\sum_{n=0}^{\infty}\,\,\sum_{W=\left(w_{1},\ldots,w_{n}\right)\in{\bf Word}(n)}X^{W}\cdot$
$\qquad\qquad\qquad\cdot\left\{\sum_{I=\left(i_{1},\ldots,i_{n}\right)\in{\bf Word}(n)}\Big{[}\theta_{i_{1}}\cdots\theta_{i_{n}}\Big{]}M_{i_{1}}\left(e_{w_{1}}\right)\bullet\cdots\bullet M_{i_{n}}\left(e_{w_{n}}\right)\right\}.$ (17)
The first equality follows from (14), the second is deduced from the definition of ${\widetilde{M}}$ (with dummy indices $(i_{1},\ldots,i_{n})$ introduced), and the final one holds by a direct rearrangement.
Now we are interested in computing the derivative of $\Phi_{X}({M_{(\theta)}})$ with respect to $\theta^{J}$ evaluated at $\theta=0\in\mathbb{R}^{d}$, where $J=\left(j_{1}\ldots,j_{L}\right)\in{\bf Word}(L)$ is a multi-index of valency $|J|=L$. That is, we shall find the matrix
$\frac{\partial^{|J|}}{\partial\theta^{J}}\Phi_{X}\left({M_{(\theta)}}\right)\bigg{|}_{\theta=0}\,\in{\mathfrak{gl}(n,\mathbb{C})}.$ (18)
The indices $j_{1},\ldots,j_{L}$ may have repetitions. Notice that for $X\in\widehat{{\mathscr{E}}}$, (18) is always well defined, since $\frac{\partial^{|J|}}{\partial\theta^{J}}\Phi_{X}\left({M_{(\theta)}}\right)$ is infinitely differentiable in $\theta$ for $|\theta|$ sufficiently small, i.e., when the operator norm of ${M_{(\theta)}}$ is less than $\rho(X)$.
To this end, notice that the last line of (3) lies in
${\mathfrak{gl}(n,\mathbb{C})}\otimes\mathbb{R}[\theta_{1},\ldots,\theta_{d}];$
i.e., it is a matrix-valued $d$-variate polynomial in $\theta_{1},\ldots,\theta_{d}$. Thus, the nontrivial contributions to the quantity in (18) arise only from those terms in the right-most side of (3) with
  • words $W=\left(w_{1},\ldots,w_{n}\right)$ of length $n=L$; and
  • multi-indices $I:=(i_{1},\ldots,i_{n})\in{\{1,\ldots,d\}}^{L}$ which differ from $J=\left(j_{1}\ldots,j_{L}\right)$ only by a permutation; i.e., there exists $\sigma\in\Sigma_{n}$ such that $I=\sigma(J)$, namely that $j_{\sigma(\nu)}=i_{\nu}$ for each $\nu\in{\{1,\ldots,n\}}$.
Observe also that for any $I=\sigma(J)$ it holds that
$\frac{\partial^{|J|}}{\partial\theta^{J}}\bigg{|}_{\theta=0}\theta_{i_{1}}\cdots\theta_{i_{n}}={\mathcal{C}_{J}},$
where ${\mathcal{C}_{J}}$ is the combinatorial constant
${\mathcal{C}_{J}}:=\prod_{\alpha=1}^{|J|}\left[r^{J}\left(j_{\alpha}\right)!\right]=\prod_{i=1}^{d}\left[r^{J}\left(i\right)!\right].$ (19)
Here $r^{J}(i)$ is the multiplicity of $i$ in the word $J$ as before, and the convention $0!=1$ is adopted. Note too that ${\mathcal{C}_{J}}\equiv\mathcal{C}_{I}$ whenever $I\sim J$.
Summarising the above arguments, we obtain for any multi-index/word $J$ the following equality in ${\mathfrak{gl}(k,\mathbb{C})}$:
$\frac{1}{{\mathcal{C}_{J}}}\frac{\partial^{|J|}\Phi_{X}({M_{(\theta)}})}{\partial\theta^{J}}\Bigg{|}_{\theta=0}=\sum_{{\overline{W}}\in{\bf Word}({|J|})}X^{\overline{W}}\left\{\sum_{\{I\in{\bf Word}({|J|}):\,I\sim J\}}M_{i_{1}}\left(e_{{\overline{w}}_{1}}\right)\bullet\cdots\bullet M_{i_{|J|}}\left(e_{{\overline{w}}_{|J|}}\right)\right\},$ (20)
where ${\overline{W}}=\left({\overline{w}}_{1},\ldots,{\overline{w}}_{{|J|}}\right)$ and $I=\left(i_{1},\ldots,i_{|J|}\right)$. We use the dummy variable ${\overline{W}}$ here to avoid potential confusion from the subsequent developments. Note that ${\mathcal{C}_{J}}$ is independent of both ${\overline{W}}$ and $I$. The matrices $M_{1},\ldots,M_{d}\in{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)$ in (16), as well as the size of the matrices $k$, are completely arbitrary in (20).
4. Determination of tensor from the development
Now we are at the stage of proving our Main Theorem 1.8. We endeavour to explain the motivations for the choices of mathematical objects that appear in the statement of the theorem.
Proof of Theorem 1.8.
The arguments are divided into five steps below.
Step 1. Fix arbitrary $n\in\mathbb{N}$ and $W=\left(w_{1},\ldots,w_{n}\right)\in{\bf Word}(n)$. We aim to solve for the coefficient $X^{W}\in\mathbb{R}$ in front of $e_{W}$; see (14) for the notations.
In view of the comments at the end of §3, we have the freedom of choosing $\{M_{i}\}_{1\leq i\leq d}$ in (16) that depends on $W$ and allows us to determine $X^{W}$ from (20). We shall designate such choices of $M_{i}$ as ${{\mathcal{M}}[W]}_{i}$, in order to emphasise that they depend only on $W$. For hyperparameters
$\theta=(\theta_{1},\ldots,\theta_{d})\in\mathbb{R}^{d}$
we write
${{{\mathcal{M}}[W]}_{(\theta)}}=\sum_{i=1}^{d}\theta_{i}{{\mathcal{M}}[W]}_{i}.$
Here and hereafter, we require the size of the matrices to satisfy $k\geq n+1$.
Step 2. Our starting point is the formula (20) obtained in §3 above. Let us rewrite the inner summands thereof by
$\left\langle\left\langle I,{\overline{W}}\right\rangle\right\rangle_{M}:=M_{i_{1}}\left(e_{{\overline{w}}_{1}}\right)\bullet\cdots\bullet M_{i_{n}}\left(e_{{\overline{w}}_{n}}\right)\in{\mathfrak{gl}(k,\mathbb{C})}$ (21)
for ${\overline{W}}=\left({\overline{w}}_{1},\ldots,{\overline{w}}_{{n}}\right)$ and $I=\left(i_{1},\ldots,i_{n}\right)$. We then recast (20) with $W$ in place of $J$ as follows:
$\frac{1}{{\mathcal{C}_{W}}}\frac{\partial^{n}\Phi_{X}\left({M_{(\theta)}}\right)}{\partial\theta^{W}}\Bigg{|}_{\theta=0}=\sum_{\left\{\left({\overline{W}},I\right)\in{\bf Word}(n)\times{\bf Word}(n):\,I\sim W\right\}}X^{\overline{W}}\left\langle\left\langle I,{\overline{W}}\right\rangle\right\rangle_{M}.$ (22)
Notice that the comments ensuing (18) explains the necessity of the assumption $X\in\widehat{{\mathscr{E}}}$. Also, by virtue of (20), nonzero contributions to the summation on the right-hand side of (22) arises only from the pairs $(I,{\overline{W}})$ such that $I\sim W$.
Step 3. Next let us motivate our choice for $\left\{M_{i}={{\mathcal{M}}[W]}_{i}\right\}_{1\leq i\leq d}$. The essential property required for $\left\{{{\mathcal{M}}[W]}_{i}\right\}_{1\leq i\leq d}$ is the following:
The (1,n+1)-entry of $\text{The $(1,n+1)$-entry of }\left\langle\left\langle I,{\overline{W}}\right\rangle\right\rangle_{{{\mathcal{M}}[W]}}=\begin{cases}1\qquad\text{if }I=W={\overline{W}},\\ 0\qquad\text{if otherwise},\end{cases}$
[…]
Indeed, assuming (4), we find that all the terms on the right-hand side of (22) have trivial $(1,n+1)$-entry unless $I=W={\overline{W}}$. Recall the definition of ${\mathcal{C}_{W}}$ from (19); we then have
$X^{W}=\frac{1}{\prod_{i=1}^{d}\left[r^{W}\left(i\right)!\right]}\times\left\{\text{the $(1,n+1)$-entry of }\frac{\partial^{n}\Phi_{X}\left({{{\mathcal{M}}[W]}_{(\theta)}}\right)}{\partial\theta^{W}}\Bigg{|}_{\theta=0}\right\},$ (24)
which is precisely the formula (13) to be verified.
Step 4. We specify $\left\{{{\mathcal{M}}[W]}_{i}\right\}_{1\leq i\leq d}$ as in the statement of the theorem, namely that
${{\mathcal{M}}[W]}_{i}(e_{j}):=\begin{cases}\delta_{ij}\left\{\sum_{t=1}^{r^{W}(i)}\left(E^{p^{W}(i|t)}_{p^{W}(i|t)+1}-E^{p^{W}(i|t)+1}_{p^{W}(i|t)}\right)\right\}\quad\text{ if }r^{W}(i)\neq 0,\\ 0\qquad\text{ if }r^{W}(i)=0.\end{cases}$ (25)
We reiterate that these matrices lie in ${\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{so}}(k,\mathbb{R})\subset{\mathfrak{su}}(k)\right)$, where $k\geq n+1$ is arbitrary.
Recall here that $E^{a}_{b}\in{\mathfrak{gl}(k,\mathbb{C})}$ are the elementary matrices. The multiplicity function $r^{W}(i)$ and position function $p^{W}(i|t)$ are as in Definition 2.1. Note too that ${{\mathcal{M}}[W]}_{i}(e_{j})$ have only nonzero entries (which are $\pm$) on the super- and sub-diagonal, i.e., at $(q,q\pm 1)$-coordinates.
Step 5. Now let us check that the choice of matrices in (25) indeed satisfies the condition (4). All we need to do is to evaluate the $(1,n+1)$-entry of
$\left\langle\left\langle I,{\overline{W}}\right\rangle\right\rangle_{{{\mathcal{M}}[W]}}={{\mathcal{M}}[W]}_{i_{1}}\left(e_{{\overline{w}}_{1}}\right)\bullet\cdots\bullet{{\mathcal{M}}[W]}_{i_{n}}\left(e_{{\overline{w}}_{n}}\right)\qquad\text{where }I\sim W.$
We refer to (21) for relevant notations, and to (22) and the ensuing remark in Step 2 for the reason of restricting to $I\sim W$.
This is completed by the two cases below.
Case 1: ${\overline{W}}\neq I$. There exists some $\ell\in{\{1,\ldots,n\}}$ such that $i_{\ell}\neq{\overline{w}}_{\ell}$. Thus
${{\mathcal{M}}[W]}_{i_{\ell}}\left(e_{{\overline{w}}_{\ell}}\right)=0,$
and hence $\left\langle\left\langle I,{\overline{W}}\right\rangle\right\rangle_{{{\mathcal{M}}[W]}}$ equals the zero matrix.
Case 2: $I={\overline{W}}$. In this case we are interested in the term
$\left\langle\left\langle I,I\right\rangle\right\rangle_{{{\mathcal{M}}[W]}}={{\mathcal{M}}[W]}_{i_{1}}\left(e_{i_{1}}\right)\bullet\cdots\bullet{{\mathcal{M}}[W]}_{i_{n}}\left(e_{i_{n}}\right),$ (26)
which is a product of $n$ matrices.
For each $i\in\{i_{1},\ldots,i_{n}\}$, by definition in (25) we have
${{\mathcal{M}}[W]}_{i}(e_{i}):=\sum_{t=1}^{r^{W}(i)}\left(E^{p^{W}(i|t)}_{p^{W}(i|t)+1}-E^{p^{W}(i|t)+1}_{p^{W}(i|t)}\right).$ (27)
Thus
[…]
where
either $a_{i}+1=b_{i}$ or $a_{i}=b_{i}+1$ for each $i\in{\{1,\ldots,n\}}$.
Now, thanks to the identity (15) for elementary matrices, we have
$E^{a_{1}}_{b_{1}}\bullet\cdots\bullet E^{a_{n}}_{b_{n}}=\delta_{b_{1}}^{a_{2}}\cdots\delta_{b_{n-1}}^{a_{n}}E^{a_{1}}_{b_{n}}.$
If this term contributes nontrivially to the $(1,n+1)$-entry of $\left\langle\left\langle I,I\right\rangle\right\rangle_{{{\mathcal{M}}[W]}}$, then one must have
$a_{1}=1,\qquad b_{n}=n+1,\qquad\text{ and }a_{\ell+1}=b_{\ell}\text{ for each }\ell\in\{1,\ldots,n-1\}.$
This together with the triangle inequality and that $|a_{i}-b_{i}|=1$ for each $i\in{\{1,\ldots,n\}}$ (due to our choice of $a_{i}$, $b_{i}$) yields that
$n$ $=b_{n}-a_{1}$
$\leq|b_{n}-a_{n}|+|a_{n}-b_{n-1}|+|b_{n-1}-a_{n-1}|$
$\qquad\qquad+|a_{n-1}-b_{n-2}|+\ldots+|a_{2}-b_{1}|+|b_{1}-a_{1}|$
$=\sum_{i=1}^{n}|b_{i}-a_{i}|=n,$
so equalities must hold everywhere. But this is only possible when
$(a_{1},\ldots,a_{n})=(1,\ldots,n-1)\qquad\text{and}\qquad(b_{1},\ldots,b_{n})=(2,\ldots,n).$
Therefore, one infers from (28) and the previous paragraph that the only nontrivial contribution to the $(1,n+1)$-entry of $\left\langle\left\langle I,I\right\rangle\right\rangle_{{{\mathcal{M}}[W]}}$ comes from a single term of the form
$\lambda E^{1}_{2}\bullet E^{2}_{3}\bullet\cdots\bullet E_{n}^{n-1}\qquad\text{for some $\lambda\in\mathbb{R}$.}$ (29)
Recall the expression (26) for $\left\langle\left\langle I,I\right\rangle\right\rangle_{{{\mathcal{M}}[W]}}$; each term on the right-hand side of which is expressed via (27). The term in (29) arises only if the following condition holds:
For each {1,,n}, there is t{1,,rW(i)} such that $\text{For each $\ell\in{\{1,\ldots,n\}}$, there is $t\in\left\{1,\ldots,r^{W}(i_{\ell})\right\}$ such that $p^{W}\left(i_{\ell}|t\right)=\ell$}.$
In view of Definition 2.1, this condition is tantamount to the following:
For each {1,,n}, the index i must occur at location  in the word $\text{For each $\ell\in{\{1,\ldots,n\}}$, the index $i_{\ell}$ must occur at location $\ell$ in the word $W$}.$
This clearly implies $W=(i_{1},\ldots,i_{n})=I$.
Also, by an inspection of (26) and (27), the coefficient $\lambda$ in (29) must be 1. Thus we have verified (4), which completes the proof of Theorem 1.8. ∎
Remark 4.1.
A careful examination reveals that the term $-E^{p^{W}(i|t)+1}_{p^{W}(i|t)}$ in our expression (25) for ${{\mathcal{M}}[W]}_{i}$ is somewhat immaterial to the proof of Theorem 1.8. We include it only to ensure that ${{\mathcal{M}}[W]}_{i}$ takes value in the space ${\mathfrak{so}}(k,\mathbb{R})$ of anti-symmetric matrices. This leaves open the possibility for adapting our proof to Cartan developments into more general Lie groups.
The choice of ${{\mathcal{M}}[W]}_{i}$ is non-unique. Indeed, if we allow ${{\mathcal{M}}[W]}_{i}$ to take values in ${\mathfrak{su}}(k)$ rather than restricting to ${\mathfrak{so}}(k,\mathbb{R})$, we may choose for instance ${\sqrt{-1}}E^{p^{W}(i|t)}_{p^{W}(i|t)}+E^{p^{W}(i|t)}_{p^{W}(i|t)+1}-E^{p^{W}(i|t)+1}_{p^{W}(i|t)}$ in lieu of $E^{p^{W}(i|t)}_{p^{W}(i|t)+1}-E^{p^{W}(i|t)+1}_{p^{W}(i|t)}$ in (25).
5. Determination of laws of stochastic processes
We now discuss the implications of our Main Theorem 1.8 on the problem of determining measures of stochastic processes from the expected development. Indeed, Theorem 1.8 identifies a new family of sparse matrix-valued linear mappings distinguishing the measures associated to path signatures, thus improving the classical result [7, Theorem 4.8] by Chevyrev–Lyons.
5.1. Separation of points
Recall from (1) the space of tridiagonal anti-symmetric matrices:
$\mathfrak{t}(k,\mathbb{R})$ :=$:=\bigg{\{}M\in\mathbb{R}^{k\times k}:\,\text{$M+M^{\top}=0$, andthe $(i,j)$-entry of $M$}$
$\qquad\qquad\qquad\qquad\qquad\text{is possibly nonzero only when$|i-j|=1$}\bigg{\}}.$
Also recall Corollary 1.9 (reproduced below):
Let $X\in\widehat{{\mathscr{E}}}$ be such that $\pi_{n}(X)\neq 0$ for some $n\in\mathbb{N}$. Then, for any $k\geq n+1$ one can find $M\in\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$ such that $\Phi_{X}(M)\neq 0$.
A direct consequence is as follows (the development map $\mathscr{D}_{\gamma}$ is defined in (5)):
Corollary 5.1.
Let $\gamma_{1}$, $\gamma_{2}\in\mathcal{V}_{p}\left([0,T],\mathbb{R}^{d}\right)$ be paths on $\mathbb{R}^{d}$ of bounded $p$-variations; $21\leq p<$ 2. Suppose that their signatures differ at the $n^{\text{th}}$ level, namely that $\pi_{n}[S(\gamma_{1})]\neq\pi_{n}[S(\gamma_{2})]$. Then, for any $k\geq n+1$ one can find $M\in\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$ such that $\mathscr{D}_{\gamma_{1}}(M)\neq\mathscr{D}_{\gamma_{2}}(M)$.
Chevyrev–Lyons [7, Theorem 4.8] asserts that for any $X\in{\mathscr{E}}^{\infty}$ (which, by definition, is a subset of $\widehat{{\mathscr{E}}}$) with $\pi_{n}(X)\neq 0$, one can find $M\in{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{sp}(m)\right)$ for some $m\geq\max\left\{2,\frac{n}{3}\right\}$ such that $\Phi_{X}(M)\neq 0$. In contrast, Corollary 5.1 establishes that linear mappings taking values in the smaller matrix algebra $\mathfrak{t}(k,\mathbb{R})$ separate points over a larger subset of the tensor algebra containing the expected signatures.
5.2. Uniqueness of restricted path characteristic function
Motivated by Corollary 5.1, let us consider the restriction of the development map $\mathscr{D}_{\gamma}$ to the domain $\mathcal{L}\left(\mathbb{R}^{d};\bigoplus_{k=0}^{\infty}\mathfrak{t}(k,\mathbb{R})\right)$. This serves as the foundation for our novel applications of $\mathscr{D}_{\gamma}$ to hypothesis testing in §6 below.
Definition 5.2.
Let $21\leq p<$ 2 and $\gamma\in\mathcal{V}_{p}\left([0,T],\mathbb{R}^{d}\right)$. Let $\mathscr{D}_{\gamma}:\mathcal{L}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)\to G$ be the unitary development, namely that the Cartan development defined in (5) with $\mathfrak{g}=\bigoplus_{k=0}^{\infty}\mathfrak{su}(k)$. We call the restriction of $\mathscr{D}_{\gamma}$ to $\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}\right)$ the restricted unitary feature of $\gamma$. Its expectation is called the restricted path characteristic function (restricted PCF) of $\gamma$.
We establish the uniqueness of restricted PCF under the condition for the decay rate of expected signature by applying the uniqueness of expected signature ([7, Proposition 6.1] ):
Theorem 5.3.
Let $\gamma_{1},\gamma_{2}$ be $\mathcal{V}_{1}\left([0,T],\mathbb{R}^{d}\right)$-valued stochastic processes whose expected signatures have infinite radii of convergence. $S(\gamma_{1}){\buildrel d\over{=}}S(\gamma_{2})$ if and only if they have the same restricted PCF:
$\mathbb{E}\left[\mathcal{D}_{\gamma_{1}}(M)\right]=\mathbb{E}\left[\mathcal{D}_{\gamma_{2}}(M)\right]\qquad\text{for any }M\in{\mathcal{L}}\left(\mathbb{R}^{d};\bigoplus_{k=0}^{\infty}\mathfrak{t}(k,\mathbb{R})\right).$ (30)
Proof.
The expected signature has infinite radius of convergence ([7]), so it determines the law of signature. By Theorem 1.8, it is determined in turn by the restricted PCF. ∎
As an illustrative example, we apply our Main Theorem 1.8 to recover the expected signature of the $d$-dimensional Brownian motion (BM) from its expected unitary development.
Example 5.4 (Recovering the expected signature of BM).
Denote the BM on time interval $[0,T]$ as $B_{[0,T]}=\left(B^{1}_{[0,T]},B^{2}_{[0,T]},\cdots,B^{d}_{[0,T]}\right)$. The expected signature of $B_{[0,T]}$ is given in [22]:
$\mathbb{E}\left[S(B_{[0,T]})\right]=\exp\left(\frac{T}{2}\sum_{i=1}^{d}e_{i}\otimes e_{i}\right).$ (31)
It is defined via iterated integrals in the sense of Stratonovich, a.k.a. the canonical Brownian rough path [17]. Then, by definition, the generating function of $B_{[0,T]}$ is
$\Phi_{B}(M_{(\theta)}):=\widetilde{M}_{(\theta)}\Big{[}\mathbb{E}(S(B_{[0,T]}))\Big{]}=\exp\left\{\frac{T}{2}\sum_{i=1}^{d}\left[M_{(\theta)}(e_{i})\right]^{2}\right\}.$
Let $W=(w_{1}w_{2}\cdots w_{L})$ be any fixed word of length $L$. We aim at determining the $W$-component of BM’s expected signature, $\mathbb{E}[S(B_{[0,T]})]_{W}$. For notational simplicity, write $M_{i}^{W}(e_{i})$ as $M_{i}$ for $i\in\{1,2,\ldots,d\}$. Differentiate the generating function $\Phi_{B}$ to obtain that
$\frac{\partial^{L}\big{[}\Phi_{B}(M_{(\theta)})\big{]}}{\partial\theta_{w_{1}}\partial\theta_{w_{2}}\cdots\partial\theta_{w_{L}}}\bigg{|}_{\theta=0}=\begin{cases}0\qquad\text{if $r^{W}(i)$ is odd for some $1\leq i\leq d$},\\ \frac{\mathcal{C}_{W}}{\left(\frac{L}{2}\right)!}\left(\frac{T}{2}\right)^{\frac{L}{2}}\sum_{I\in\widetilde{\Sigma}(W)}\left(M_{i_{1}}M_{i_{2}}\cdots M_{i_{L}}\right)\qquad\text{if else},\end{cases}$ (32)
where
$\widetilde{\Sigma}(W):=\left\{I\in\widetilde{\sigma}(W):\,I_{p}=I_{p+1}\text{ for each odd $p\in\{1,\ldots,L\}$}\right\};$
$\widetilde{\sigma}(W):=\left\{\left(w_{\tau(1)}w_{\tau(2)}\cdots w_{\tau(L)}\right):\,\tau\in\Sigma_{L}\right\}.$
For example, $\widetilde{\Sigma}((1212))=\{(1122),(2211)\}$. As a result, the expected signature computed via Theorem 1.8 is
$\widehat{\mathbb{E}}\left[S\left(B_{[0,T]}\right)\right]_{W}=\begin{cases}\frac{1}{\left(\frac{L}{2}\right)!}\left(\frac{T}{2}\right)^{\frac{L}{2}},&\text{ if }r^{W}(i)\text{ is even and }W\in\widetilde{\Sigma}(W)\text{ for each }i\in\{1,2,\ldots,d\},\\ 0,&\text{ otherwise},\end{cases}$
which coincides with the known formula (31).
5.3. Uniqueness of orthogonal path characteristic function
As a variant of the restricted PCF in Definition 5.2, we introduce the following:
Definition 5.5.
Let $21\leq p<$ 2 and $\gamma\in\mathcal{V}_{p}\left([0,T],\mathbb{R}^{d}\right)$. Let $\mathscr{D}_{\gamma}:\mathcal{L}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)\to G$ be the unitary development, namely that the Cartan development defined in (5) with $\mathfrak{g}=\bigoplus_{k=0}^{\infty}\mathfrak{su}(k)$. We call the restriction of $\mathscr{D}_{\gamma}$ to $\mathcal{L}\left(\mathbb{R}^{d};\bigoplus_{k=0}^{\infty}\mathfrak{so}(k,\mathbb{R})\right)$ the orthogonal feature of $\gamma$. Its expectation is called the orthogonal path characteristic function (orthogonal PCF) of $\gamma$.
The special orthogonal matrix algebra ${\mathfrak{so}}(k,\mathbb{R})$ consists of anti-symmetric real matrices. Note for each $k\in\mathbb{N}$ that
$\mathfrak{t}(k,\mathbb{R})\subset{\mathfrak{so}}(k,\mathbb{R})\subset{\mathfrak{su}}(k).$
Without any additional regularity assumptions on the expected signature, we have:
Theorem 5.6.
Let $\gamma_{1},\gamma_{2}$ be $\mathcal{V}_{1}\left([0,T],\mathbb{R}^{d}\right)$-valued stochastic processes. Then $S(\gamma_{1}){\buildrel d\over{=}}S(\gamma_{2})$ if and only if their OPCF are the same, i.e.,
$\mathbb{E}\left[\mathcal{D}_{\gamma_{1}}(M)\right]=\mathbb{E}\left[\mathcal{D}_{\gamma_{2}}(M)\right]\qquad\text{for any }M\in{\mathcal{L}}\left(\mathbb{R}^{d};\bigoplus_{k=0}^{\infty}\mathfrak{so}(k,\mathbb{R})\right).$ (33)
Proof.
By Corollary 1.9 and $\mathfrak{t}(k,\mathbb{R})\subset{\mathfrak{so}}(k,\mathbb{R})$, the orthogonal feature separates points over tensors of positive radii of convergence. We claim that $L\circ\mathcal{D}_{\gamma}(M)$ where $L$ is a linear functional and $M\in{\mathcal{L}}\left(\mathbb{R}^{d};\bigoplus_{k=1}^{\infty}{\mathfrak{so}}(k)\right)$ forms an algebra; i.e., for any $(L_{1},M_{1})$ and $(L_{2},M_{2})$ there exists $(L_{3},M_{3})$ such that $\Big{(}L_{1}\circ\widetilde{M_{1}}\Big{)}\bullet\Big{(}L_{2}\circ\widetilde{M_{2}}\Big{)}=L_{3}\circ\widetilde{M_{3}}$. Say $M_{i}\in{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{so}}(k_{i})\right)$ for $i\in\{1,2,3\}$.
To see this claim, simply note that
$\widetilde{M_{1}}(S(\gamma))\otimes\widetilde{M_{2}}(S(\gamma))=\widetilde{M_{3}}(S(\gamma))$
for
$M_{3}=M_{1}\otimes I_{k_{2}}+I_{k_{1}}\otimes M_{2},$
and that $\mathcal{D}_{\gamma}(M)$ equals the constant $1$ for $k=1$ and $M=0$. Thus, by Stone–Weierstrass [7, Lemma 4.2], the characteristic function defined in (12) with $M$ restricted to $\bigoplus_{k=1}^{\infty}{\mathfrak{so}}(k)$ determines the Borel measures on $\mathcal{G}\left(\mathbb{R}^{d}\right)$. ∎
6. Numerical experiments on hypothesis testing
6.1. A new distance for stochastic processes
According to Theorem 5.3, under the regularity condition therein, the restricted PCF of a stochastic process determines its law (and vice versa). This motivates the introduction of a new distance function for probability measures on the path space, termed as the restricted path characteristic function distance (RPCFD), by restricting the unitary development/feature to the sparse subspace $\mathfrak{t}(k,\mathbb{R})$. Similarly, one can define a orthogonal PCF-based distance function. These natural variants of the path characteristic function distance (PCFD) proposed in [15] demonstrate significant computational benefits due to the sparsity of $\bigoplus_{k=0}^{\infty}\mathfrak{t}(k;\mathbb{R})$ and $\bigoplus_{k=0}^{\infty}\mathfrak{so}(k;\mathbb{R})$.
Denote the Hilbert–Schmidt norm on $\mathbb{C}^{m\times m}=\mathfrak{gl}(m,\mathbb{C})$ as $\|\cdot\|_{\rm HS}$:
$||A||_{\rm HS}:=\sqrt{\operatorname{tr}(AA^{\ast})}.$
In the sequel we shall focus on the methodology of learning the RPCFD.
In view of Hambly–Lyons [10] and Theorem 5.3, the restricted unitary feature of a path of bounded variation only determines the path modulo tree-like equivalence. However, tree-like equivalence is mod out by lifting to time-augmented paths:
$\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right):=\left\{\tilde{X}:\,\tilde{X}(t)=(t,X(t))\text{ for }X\in\mathcal{V}_{p}\left([0,T],\mathbb{R}^{d}\right)\right\}.$
Here $\mathcal{V}_{p}\left([0,T],\mathbb{R}^{d}\right)$ denotes the space of bounded $p$-variation paths into $\mathbb{R}^{d}$. We also set
$\mathcal{P}_{0}\left({\tilde{\mathcal{V}}_{1}([0,T],\mathbb{R}^{d})}\right):=\Bigg{\{}\text{Borel probability measures on $\tilde{\mathcal{V}}_{1}\left([0,T],\mathbb{R}^{d}\right)$ }$
$\qquad\qquad\text{under which the corresponding expected signature has ${\rm ROC}=\infty$}\Bigg{\}}.$ (34)
Then we introduce:
Definition 6.1.
Fix $k\geq 1$ and any Borel probability measure $\mathcal{M}$ on $\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$ with full support. For $\mu,\nu\in\mathcal{P}_{0}\left({\tilde{\mathcal{V}}_{1}\left([0,T],\mathbb{R}^{d}\right)}\right)$, their restricted path characteristic function distance (RPCFD) with respect to $\mathcal{M}$ is defined as
${\rm RPCFD}_{\mathcal{M}}^{2}(\mu,\nu):=\mathbb{E}_{M\sim\mathcal{M}}\left[\big{|}\big{|}\Phi_{\mu}(M)-\Phi_{\nu}(M)\big{|}\big{|}_{\rm HS}^{2}\right],$ (35)
with the characteristic functions $\Phi_{\mu}(M):=\mathbb{E}_{X\sim\mu}\left[\widetilde{M}(S(X))\right]$ and $\Phi_{\nu}(M):=\mathbb{E}_{Y\sim\nu}\left[\widetilde{M}(S(Y))\right]$.
As $\mathfrak{t}(k,\mathbb{R})\subset{\mathfrak{su}}(k)$, RPCFD can be viewed as a restricted version of PCFD that inherits various nice properties from PCFD, including the uniform boundedness and differentiability with respect to model parameters. Moreover, since
$\dim(\mathfrak{t}(k,\mathbb{R}))=k-1<\dim({\mathfrak{su}}(k))=k^{2},$
RPCFD offers a significant dimension reduction benefit (i.e., in need of fewer parameters) over PCFD while retaining the characteristic property.
Remark 6.2.
Likewise, based on Theorem 5.6 one may define the orthogonal path characteristic function distance (OPCFD) for stochastic processes by restricting the domain to $\bigoplus_{k=0}^{\infty}{\mathfrak{so}}(k)$. As $dim({\mathfrak{so}}(k))=k(k-1)/2$, RPCFD reduces the dimension more significantly than OPCFD, so we focus more on RPCFD. Nonetheless, most of the analysis and numerical methods below also apply to OPCFD. See Appendix B for numerical experiments of OPCFD.
To enhance the discriminative power of RPCFD on given path distributions, we propose the following trainable distance:
$\sup_{\mathcal{M}\in\mathcal{P}(\mathcal{L}(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})))}\text{RPCFD}^{2}_{\mathcal{M}}(\mu,\nu).$ (36)
In practice, we approximate $\mathcal{M}$ by an empirical measure:
$\mathcal{M}\approx\frac{1}{K}\sum_{i=1}^{K}\delta_{M_{i}},$ (37)
where $M_{i}\in\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$, $i\in\{1,\cdots,K\}$, for some $K\geq 1$.
Note that the linear mappings $\{M_{i}\}_{1\leq i\leq K}$ can be fully parameterised. Therefore, the task of learning the optimal distribution $\mathcal{M}$ in (36) is transformed to learning the optimal parameters in the linear mappings $\{M_{i}\}_{1\leq i\leq K}$ which attain the supremum of the “empirical RPCFD”:
$\sup_{\{M_{i}\}_{1\leq i\leq K}}\frac{1}{K}\sum_{i=1}^{K}\text{RPCFD}^{2}_{\delta_{M_{i}}}(\mu,\nu):=\sup_{\{M_{i}\}_{1\leq i\leq K}}\frac{1}{K}\sum_{i=1}^{K}||\Phi_{\mu}(M_{i})-\Phi_{\nu}(M_{i})||_{\rm HS}^{2}.$ (38)
6.2. Numerical experiments on hypothesis testing
We apply the trainable empirical RPCFD in (38) to the task of hypothesis testing on laws of stochastic processes. The empirical RPCFD acts as the statistic. With an abuse of notation, suppose $X$ and $Y$ are two stochastic processes in the Euclidean space. The hypotheses of our two-sample permutation test are:
[…]
where $\mathbb{P}_{X}$ and $\mathbb{P}_{Y}$ denote the laws of $X$ and $Y$, respectively.
Denoted by $\mathbf{X}=\{X_{1},X_{2},\cdots,X_{m}\}$ and $\mathbf{Y}=\{Y_{1},Y_{2},\cdots,Y_{n}\}$ two sets of i.i.d. samples drawn respectively from $\mathbb{P}_{X}$ and $\mathbb{P}_{Y}$. Given the statistic $T(\mathbf{X},\mathbf{Y})$ (i.e., the empirical RPCFD), we apply all possible the permutations $\sigma\in\Sigma_{m+n}$ to $\mathbf{Z}:=(\mathbf{X},\mathbf{Y})$ and obtain the set $\mathcal{T}$ via
$\mathcal{T}:=\bigg{\{}T(\mathbf{Z}_{\sigma(1):\sigma(m)},\mathbf{Z}_{\sigma(m+1):\sigma(m+n)})\;\left|\;\sigma\in\Sigma_{m+n}\bigg{\}}\right..$
Under $H_{0}$, one can uniformly sample from $\mathcal{T}$ to approximate the distribution of $T(\mathbb{P}_{X},\mathbb{P}_{Y})$. Then, given the significance level $\alpha$, the null hypothesis $H_{0}$ will be rejected if $T(\mathbf{X},\mathbf{Y})$ exceeds the $(1-\alpha)\%$ quantile of $\mathcal{T}$, otherwise accepted. More details on the two-sample permutation test can be found in e.g. [26, §7].
In practice, we implicitly assume that the conditions of Theorem 5.3 are satisfied for $X$ and $Y$. This assumption is validated by the fact that in practice, only discrete data points are available and interpolation methods have to be applied. Keeping this in mind, we follow the workflow of [15]. Specifically, given the training data, which is i.i.d. generated under $H_{1}$, we first optimise the parameters of the empirical RPCFD via (38). Then, we use the optimised empirical RPCFD to conduct the two-sample permutation test on another out-of-sample test dataset, whose samples are generated under either $H_{0}$ or $H_{1}$. Here we use the test power (true rejection, the higher the better) and Type-I error (false rejection, the lower the better) as the metrics of the performance of our optimised empirical RPCFD.
Following [15, Example B.12], we set $X:=\tilde{B}_{[0,T]}$ to be the time-augmented 3-dimensional Brownian Motion (BM), and $Y:=\tilde{B}^{h}_{[0,T]}$ to be the time-augmented 3-dimensional fractional Brownian Motion (fBM) with the Hurst parameter $h\in]0,1[$. Note that $\mathbb{P}_{X}\neq\mathbb{P}_{Y}$ when $h\neq 0.5$.
To benchmark, we consider other distance-based statistics (i.e., PCFD), as well as signature-based MMDs [26]. Detailed descriptions of the experiment hyperparameters and results can be found in Appendices A and B. For readers interested in reproducing the numerical experiments, we have made the code implementation publicly available at the GitHub repository: https://github.com/DeepIntoStreams/UnitaryDevInversion.
RPCFD vs. PCFD
We compare our proposed trainable empirical RPCFD with PCFD proposed in [15]. Recall $K$ and $k$ from §6.1. In the experiments, we set $K=8$ and $k=5$. As shown in Figure 6.1 (Upper Left), the overall test power curves against various $h$ for RPCFD and PCFD are close. In the two cases $h=0.45$ and $h=0.55$, RPCFD achieves an obviously superior test power over PCFD, indicating that restricting the linear maps to $\mathfrak{t}(k,\mathbb{R})$ effectively narrows the search space, thereby making the learning more efficient. Moreover, due to the sparsity of the linear maps in RPCFD (and that RPCFD only uses real numbers), the training time of RPCFD is also notably reduced by $22.9\%$ on average; see the upper right panel of Figure 6.1.
Figure 6.1. (Upper Left) Test power of trained RPCFD and PCFD, where the shaded area represents mean $\pm$ std over 5 runs. (Upper Right) Type-I error of trained RPCFD and PCFD. (Bottom) Training time of RPCFD and PCFD over 500 iterations. Here fix $K=8$ and $k=5$.
The sensitivity analysis of RPCFD when $|h-0.5|\in[0.01,0.04]$ is plotted in Figure 6.2 below. In general, increasing the matrix order $k$ improves the discriminative power of RPCFD. However, the larger $k$ may have the potential issues of overfitting and high computational cost. Therefore, in practice, one should choose the matrix order $k$ by hyperparameter tuning.
Figure 6.2. Sensitivity analysis of RPCFD (Left: mean test power over 5 runs; Right: mean Type-I error over 5 runs). The $y$-axis and $x$-axis represent the power of the trained RPCFD and the Hurst parameter $h$, respectively. Here fix $K=8$.
Comparison with Sig-MMDs
In Appendix B (Tables B.1 and B.2), the development-based distances RPCFD, OPCFD and PCFD outperform the signature-based MMDs consistently for varying $h$ in terms of the test power. In particular, when $h\in\{0.45,0.55\}$, the development-based distances improve the test power from $[0.15,0.17]$ to $[0.95,0.99]$. Moreover, the inference time of Sig-MMD-based methods scales quadratically with the sample size, while the dependency is linear for development-based methods; see Table B.3 for the inference time of each method across different sample sizes. For example, when sample size $m=n=200$, the inference time of signature MMDs is over 10 times longer than that of the distance-based methods.
7. Conclusion
This paper provides a new, constructive proof for solving the generalised moment problem on tensor algebra. On the theoretical side, it enables us to generalise the classical results on the unique determination of signature by unitary development [7] — both by relaxing the condition on infinite radius of convergence (ROC) to nonzero ROC, and by extending to Cartan developments into a variety of matrix Lie groups. This motivates us to propose a novel, efficient metric for probability measures on the path space, which features on the sparsity of the linear maps involved and hence brings about computational benefits in dimension reduction. We demonstrate its practical power in hypothesis testing for the law of stochastic processes. It would be worthwhile to investigate in the future its applications to synthetic time series generation.
Appendices
Appendix A Numerical Experiment Details
Permutation test
We first provide a detailed description of the algorithm for estimating test power and Type-I error in the two-sample permutation test in §6.
$\alpha\in(0,1)$: significance level; $N>0$: # of experiments; $M>0$: # of permutations; $\mathbf{X}$: samples from distribution $\mu$; $\mathbf{Y}$: samples from distribution $\nu$; $m>0$: sample size of $\mathbf{X}$; $n>0$: sample size of $\mathbf{Y}$; $T$: test statistic function; $H_{0}\in\{1,0\}$: whether the null hypothesis is true or false ($H_{0}=1$ if $\mu=\nu$, otherwise $H_{0}=0$)
$\mathbf{Z}\leftarrow$ Concatenate$(\mathbf{X},\mathbf{Y})$
[…]
$i\leftarrow 1$
while $i\leq N$ do
     $\mathcal{T}\leftarrow$ EmptyList
     $j\leftarrow 1$
     while $j\leq M$ do
         $\sigma\sim$ Permutation($\{1,2,\cdots,m+n\}$)
         $T_{\sigma}\leftarrow T(\{\mathbf{Z}_{\sigma(1)},\mathbf{Z}_{\sigma(2)},\cdots,\mathbf{Z}_{\sigma(m)}\},\{\mathbf{Z}_{\sigma(m+1)},\cdots,\mathbf{Z}_{\sigma(m+n)}\})$
         $\mathcal{T}\text{.append}(T_{\sigma})$
         $j\leftarrow j+1$
     end while
     if $T(\mathbf{X},\mathbf{Y})>(1-\alpha)\%$ quantile of $\mathcal{T}$ then
[…]
     end if
     $i\leftarrow i+1$
end while
[…]
if $H_{0}$ then
[…]
     return Type_I_error
else
[…]
     return test_power
end if
Algorithm 1 Estimating the test power/Type-I error of the permutation test
Implementation details
We next provide the full details of implementation of the numerical experiments in §6, including the choice of hyperparameters and the training procedure.
All experiments are implemented on a Tesla V100 GPU. The random seed of PyTorch and NumPy is set to be 0. When estimating the test power, 20 experiments are implemented (i.e., $N=20$). In each experiment, 200 BM and 200 fBM trajectories are sampled as the out-of-sample test dataset (i.e., the construction of $\mathbf{X}$ and $\mathbf{Y}$ with $m=n=200$), which is permuted 500 times to estimate the distribution of the permuted distance (i.e., $M=500$). The significance level of the permutation test is set to be 5% (i.e., $\alpha=0.05$).
When training each development-based distance (i.e., RPCFD, OPCFD and PCFD), 10000 BM and 10000 fBM trajectories (50 time steps on $[0,1]$) are sampled as the training set. The training utilizes PyTorch’s Adam optimiser with the $\beta$’s being $(0,0.9)$. The fine-tuned learning rate of OPCFD and PCFD is 0.05 whereas that of RPCFD is 0.5. In each of the 500 iterations of the training, a mini-batch of 1024 trajectories is randomly selected from the training set to conduct the stochastic gradient ascent to maximise the corresponding empirical distance. After training, a pool of 10000 BM and 10000 fBM trajectories is independently sampled and used to conduct the permutation test.
For the signature MMDs, we consider the linear and RBF kernels (see [26, §7.1] for more details). For the RBF kernel, the fine-tuned $2\sigma^{2}$ is set to be 0.1.
Appendix B Supplementary numerical results
OPCFD vs. PCFD
As shown in Figure B.1, OPCFD has a close performance as PCFD while achieving an average time reduction of about 31.0%.
Figure B.1. (Upper Left) Test power of trained OPCFD and PCFD, where the shaded area represents mean $\pm$ std over 5 runs. (Upper Right) Type-I error of trained OPCFD and PCFD. (Bottom) Training time of OPCFD and PCFD over 500 iterations. Here fix $K=8$ and $k=5$.
Summarising tables
Comprehensive tables for summarising the test power, Type-I error and time complexity of the distances/models involved in this paper are presented below.
Developments Signature MMDs
$h$ RPCFD OPCFD PCFD Linear RBF
0.2 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$
0.25 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$
0.3 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $0.9\pm 0.04$ $\textbf{1}\pm 0$
0.35 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $0.35\pm 0.07$ $\textbf{1}\pm 0$
0.4 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $0.09\pm 0.06$ $0.97\pm 0.03$
0.425 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $0.1\pm 0.05$ $0.69\pm 0.11$
0.45 $\textbf{0.99}\pm 0.02$ $0.92\pm 0.07$ $0.74\pm 0.17$ $0.04\pm 0.04$ $0.15\pm 0.05$
0.475 $0.06\pm 0.02$ $0.12\pm 0.09$ $\textbf{0.14}\pm 0.05$ $0.01\pm 0.02$ $0.04\pm 0.02$
0.525 $0.08\pm 0.02$ $0.1\pm 0.04$ $\textbf{0.16}\pm 0.05$ $0.05\pm 0.02$ $0.07\pm 0.04$
0.55 $\textbf{0.95}\pm 0.03$ $0.82\pm 0.16$ $0.85\pm 0.16$ $0.13\pm 0.05$ $0.17\pm 0.04$
0.575 $\textbf{1}\pm 0$ $0.99\pm 0.02$ $0.97\pm 0.06$ $0.07\pm 0.02$ $0.5\pm 0.10$
0.6 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $0.05\pm 0.03$ $0.75\pm 0.05$
0.65 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $0.14\pm 0.05$ $\textbf{1}\pm 0$
0.7 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $0.23\pm 0.05$ $\textbf{1}\pm 0$
0.75 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $0.28\pm 0.07$ $\textbf{1}\pm 0$
0.8 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $0.46\pm 0.07$ $\textbf{1}\pm 0$
Table B.1. Test power of the distances when $h\neq 0.5$ in the form of mean $\pm$ std over 5 runs. For RPCFD, OPCFD and PCFD, fix $K=8$ and $k=5$. For the RBF signature MMD, fix $0.12\sigma^{2}=$ 0.1.
Developments Signature MMDs
$h$ RPCFD OPCFD PCFD Linear RBF
0.2 $0.07\pm 0.05$ $0.08\pm 0.07$ $0.04\pm 0.05$ $0.04\pm 0.09$ $0.04\pm 0.04$
0.25 $0.03\pm 0.02$ $0.02\pm 0.04$ $0.05\pm 0.05$ $0.03\pm 0.03$ $0.04\pm 0.04$
0.3 $0.05\pm 0$ $0.02\pm 0.02$ $0.04\pm 0.04$ $0.04\pm 0.04$ $0.01\pm 0.02$
0.35 $0.04\pm 0.04$ $0.05\pm 0.03$ $0.01\pm 0.02$ $0.03\pm 0.04$ $0.07\pm 0.07$
0.4 $0.04\pm 0.04$ $0.04\pm 0.04$ $0.04\pm 0.02$ $0.06\pm 0.05$ $0.04\pm 0.04$
0.425 $0.08\pm 0.07$ $0.05\pm 0.04$ $0.01\pm 0.02$ $0.03\pm 0.03$ $0.04\pm 0.04$
0.45 $0.08\pm 0.02$ $0.08\pm 0.09$ $0.09\pm 0.07$ $0.05\pm 0.04$ $0.02\pm 0.03$
0.475 $0.02\pm 0.04$ $0.07\pm 0.07$ $0.05\pm 0.03$ $0.05\pm 0.04$ $0.07\pm 0.06$
0.525 $0.09\pm 0.07$ $0.02\pm 0.02$ $0.04\pm 0.06$ $0.03\pm 0.03$ $0.04\pm 0.02$
0.55 $0.03\pm 0.04$ $0.05\pm 0.08$ $0.06\pm 0.05$ $0.07\pm 0.06$ $0.05\pm 0.05$
0.575 $0.04\pm 0.04$ $0.04\pm 0.04$ $0.1\pm 0.03$ $0.02\pm 0.03$ $0.06\pm 0.07$
0.6 $0.06\pm 0.04$ $0.04\pm 0.04$ $0.04\pm 0.04$ $0.05\pm 0.04$ $0.05\pm 0.04$
0.65 $0.02\pm 0.02$ $0.07\pm 0.02$ $0.1\pm 0.07$ $0.07\pm 0.06$ $0.03\pm 0.04$
0.7 $0.08\pm 0.05$ $0.04\pm 0.04$ $0.06\pm 0.05$ $0.05\pm 0.05$ $0.03\pm 0.04$
0.75 $0.05\pm 0.03$ $0.07\pm 0.07$ $0.05\pm 0.04$ $0.06\pm 0.07$ $0.04\pm 0.07$
0.8 $0.01\pm 0.02$ $0.05\pm 0.04$ $0.06\pm 0.02$ $0.07\pm 0.07$ $0.04\pm 0.04$
Table B.2. Type-I error of the distances when $h\neq 0.5$ in the form of mean $\pm$ std over 5 runs. For RPCFD, OPCFD and PCFD, fix $K=8$ and $k=5$. For the RBF signature MMD, fix $0.12\sigma^{2}=$ 0.1.
Developments Signature MMDs
RPCFD OPCFD PCFD Linear RBF
$m=n$ Inference time (seconds)
$10$ $175.35\pm 8.24$ $168.26\pm 10.34$ $167.69\pm 2.20$ $95.12\pm 0.21$ $122.9\pm 0.32$
$50$ $212.54\pm 8.78$ $204.02\pm 3.68$ $236.21\pm 5.46$ $402.69\pm 0.23$ $533.43\pm 0.33$
$100$ $333.08\pm 9.23$ $319.57\pm 3.01$ $376.69\pm 4.25$ $1329.73\pm 0.57$ $1760.18\pm 0.29$
$200$ $398.87\pm 22.3$ $400.63\pm 26.06$ $510.71\pm 20.08$ $5257.11\pm 0.98$ $6958.81\pm 1.33$
Mini-batch size Training time (seconds over 500 iterations)
1024 $92.46\pm 4.93$ $86.53\pm 3.07$ $123.71\pm 1.56$- --
Table B.3. Inference time of the permutation test across different sample sizes ($m=n$) and the training time of RPCFD, OPCFD and PCFD before conducting the permutation test. The result is in the form of mean $\pm$ std over $5$ runs. For RPCFD, OPCFD and PCFD, fix $K=8$ and $k=5$. For the RBF signature MMD, fix $0.12\sigma^{2}=$ 0.1. Here fix $h=0.45$.
Acknowledgement. We would like to thank Dr. Lei Jiang for his help with running the codes on the servers. SL and HN are deeply indebted to Professor Haojie Ren for insightful discussions on hypothesis testing.
Funding sources. The research of SL is supported by NSFC Projects 12201399, 12331008, and 12411530065, Young Elite Scientists Sponsorship Program by CAST 2023QNRC001, the National Key Research &\&& Development Program 2023YFA1010900, and the Shanghai Frontier Research Institute for Modern Analysis. HN is supported by the EPSRC under the program grant EP/S026347/1 and the Alan Turing Institute under the EPSRC grant EP/N510129/1. Both SL and HN are partially funded by the SJTU-UCL joint seed fund WH610160507/067 and the Royal Society International Exchanges 2023 grant (IEC/NSFC/233077).
Competing interests statement. All the authors declare that there is no conflict of interest.